Prove that (-a)^2 is equal to a^2. Use this theorem (-a)(-b)=ab


1
Expert's answer
2021-09-06T06:04:14-0400

The product of two negative numbers or terms is positive:


"(-a)(-b)=ab"

Then


"(-a)^2=(-a)(-a)=a(a)=a^2"

Therefore


"(-a)^2=a^2"